5(2x - 7) = 33 - 3(6x + 4)
Answer by stanbon(75887)

5(2x - 7) = 33 - 3(6x + 4)
---
10x - 35 = 33 - 18x -12
---
10x - 35 = 21 - 18x
---
28x = 56
---
x = 2
=========
Cheers,
Stan H.